There are 3 caravans each costing $58,000. The first caravan models reducing balance depreciation at a rate of 4.9% p.a. The second, flat rate depreciation at a rate of 4.3% p.a. Lastly, the third model’s unit cost depreciation at a rate of $1.50/km, where you hope to travel up to 1,700 km/year. Investigate under which circumstances you would choose each of these plans. Would your conclusions change if you were to travel much less or much more than 1,700km/year?
1
Expert's answer
2020-05-18T18:54:51-0400

The third caravan depreciation per year"=1,700*1.5=2550"

The second caravan depreciation per year:

"0.043*58,000=2494"

The first caravan depreciation per year:

"0.049*58,000=2842"

I choos the second caravan.

If I will travel less than "2494\/1.5=1662.6" km/year Iwill choos the third caravan otherwise the second.

If I hope to travel up to 1662.6 km/year or more the firs caravan I will choose if its depreciation at a rate will be less than 4.3% or cost will be less than "2494\/0.049=50,897.96"